One year ago a company sold 10-year, $1,000 par value, annual coupon bonds at a price of $950.00 per bond. At that time the market rate was 8 percent a year. Today the market rate is 8.5 percent a year. At what price are the bonds currently selling?
Answer #1

First of all we need to find coupon rate

Market price of bond = Interest *PVIFA(r%,n) + Redemption value*PVIF(r%,n)

950 = Interest*PVIFA(8%,10) + 1000*PVIF(8%,10)

950 = Interest*6.710081 + 1000*0.463193

950 = Interest*6.710081 +463.1935

486.8065 = Interest*6.710081

Interest = 72.5485$

Interest in % = 72.5485/1000 = 7.25%

Now let us find price of bond today

Market price of bond = Interest *PVIFA(r%,n) + Redemption value*PVIF(r%,n)

=72.5485*PVIFA(8.5%,9) + 1000*PVIF(8.5%,9)

=72.5485*6.119063 +1000*0.47988

=443.929+479.88

=923.809$